# FT: Unlocking the Power of Code Instruction Tuning by Simply Merging Upcycled Mixture-of-Experts

### v1

@@ A.5 Analysis on Expert Specialization
Inspired by recent works Jiang et al. (2024); Xue et al. (2024), we analyze whether each expert in MoE has different specializations in different programming languages by visualizing the routing decision of the tokens from different programming languages in the MultiPL-E benchmark (including Python). For the MultiPL-E benchmark, we collect the routing decision when conducting experiments in Section 4.3. For Python, we collect the routing decision by reruning HumanEval experiment following the same setting as Section 4.3. Following Mixtral Jiang et al. (2024), we get the visualization results from layers 0, 11, and 23 in MoE, where layer 0 and layer 23 are the first and the last layers of MoE. As is shown in Figure 4, we do not observe obvious patterns in the assignment of experts based on the programming language, which is in line with the observation reported by recent works Jiang et al. (2024); Xue et al. (2024). 1681
A.6 Training Settings for stable-code 3B
@@
We use evol-codealpaca-v1 as the training dataset. Since stable-code 3B is the base model, we upcycle a new MoE model from the base model, namely MoE. Due to limited computational resources, we construct MoE with 4 experts in one expert layer, where the top 2 experts are activated for each token, including one shared expert. Consequently, the size of MoE can be described as 43B. We use a batch size of 64 and a learning rate of 5e-5 with a linear scheduler to fine-tune MoE for 4 epochs with 500 warmup steps. Similar to FT, we obtain FT by learning mixing coefficients to merge MoE layers inside MoE as normal FFN layers, which is fine-tuned with a batch size of 64, a shared expert rate of 0.85, and a learning rate of 1e-5 with a linear schedule for 1 epoch with 125 warmup steps. Our baseline model, namely SFT, is fine-tuned for 5 (= 4 + 1) epochs with a batch size of 64, a learning rate of 5e-5, and 625 warmup steps for a fair comparison.
A.7 Training Settings for TinyLlama 1.1B
@@
Using TinyLlama 1.1B as the base model, we upcycle a new MoE model, namely MoE, from the base model. Following the setting for MoE, we construct MoE with 8 experts in one expert layer, where the top 6 experts are activated for each token, including one shared expert. As such, the number of parameters for MoE can be written as 81.1B. We use a batch size of 64 and a learning rate of 5e-5 with a linear scheduler to fine-tune MoE for 4 epochs with 240 warmup steps. To obtain FT, we learn mixing coefficients to merge MoE layers inside MoE by fine-tuning them with a batch size of 64, a shared expert rate of 0.85, and a learning rate of 2e-5 with a linear schedule for 1 epoch with 60 warmup steps. For a fair comparison, we fine-tune a baseline model SFT for 5 (= 4 + 1) epochs with a batch size of 64, a learning rate of 5e-5, and 300 warmup steps.
A.8 Theoratical Explanation Details
We consider a simplified variant of FT as below:
@@ A.8 Theoratical Explanation Details
In this simplified scenario, if we denote as the output of the model for the input , the output of this simplified MoE model for input token can be represented as . Interestingly, if we define two new dense models and , where and use the same attention layer as this MoE model while using and as the FFN layer separately, can be represented as ! Consequently, the computation process of this simplified MoE model can be viewed as ensembling the outputs of two dense models and . Meanwhile, the process of merging the upcycled MoE model back to a dense model in this simplified FT can be represented as , which is the merging of the same two dense models and .

### v2

@@ A.6 Expert Specialization Analysis
Inspired by recent works Jiang et al. (2024); Xue et al. (2024), we analyze whether each expert in MoE has different specializations in different programming languages by visualizing the routing decision of the tokens from different programming languages in the MultiPL-E benchmark (including Python). We collect the routing decision for the MultiPL-E benchmark when conducting experiments in Section 4.3. For Python, we collect the routing decision by running HumanEval experiment following the same setting used in Section 4.3. Following the analysis setting of recent work Jiang et al. (2024), we get the visualization results from layers 0, 11, and 23 in MoE, where layer 0 and layer 23 are the first and the last layers of MoE. As shown in Figure 4, we do not observe any obvious patterns in the assignment of experts based on the type of programming languages, which is in line with the findings reported by recent works Jiang et al. (2024); Xue et al. (2024).
A.7 Effect of Shared Expert Rate
The performance of the final merged dense model drops significantly when the shared expert rate grows from 0.75 to 1.00, showing that specific knowledge learned by other experts is also integral and ignoring them will lead to a significant performance drop.
| Model | HumanEval | HumenEval+ | |
| SFT | - | 61.6 | 57.3 |
| FT | 0.00 | 62.8 | 59.8 |
| 0.25 | 64.6 | 61.0 | |
| 0.50 | 65.9 | 62.8 | |
| 0.75 | 67.1 | 64.6 | |
| 1.00 | 63.4 | 60.4 |
A.8 Training Settings for stable-code 3B
@@
We use evol-codealpaca-v1 as the training dataset. Since stable-code 3B is the base model, we upcycle a new MoE model from the base model, namely MoE. We construct MoE with 4 experts in one MoE layer, where the top 2 experts are activated for each token, including one shared expert. Consequently, the size of MoE can be described as 43B. We use a batch size of 64 and a learning rate of 5e-5 with a linear scheduler to fine-tune MoE for 4 epochs with 500 warmup steps. Similar to FT, we obtain FT by learning mixing coefficients to merge MoE layers inside MoE as normal FFN layers, which is fine-tuned with a batch size of 64, a shared expert rate of 0.85, and a learning rate of 1e-5 with a linear schedule for 1 epoch with 125 warmup steps. Our baseline model, namely SFT, is fine-tuned for 5 (= 4 + 1) epochs with a batch size of 64, a learning rate of 5e-5, and 625 warmup steps for a fair comparison.
A.9 Training Settings for DeepSeek-Coder-Base 6.7B
@@
We use evol-codealpaca-v1 as the training dataset. We upcycle a new MoE model from DeepSeek-Coder-Base 6.7B, namely MoE. We construct MoE with 8 experts in one MoE layer, where the top 2 experts are activated for each token, including one shared expert. As such, MoE includes 86.7B parameters. We use a batch size of 64 and a linear scheduler to fine-tune MoE for 4 epochs with 500 warmup steps. We choose the best-performing learning rate from for MoE. Because the FFN weights of MoE are too large to fit in our GPU memory, during our merging step, we realize that one part of computation in the training of FT has to be moved to CPUs, which significantly slows down the training speed. Consequently, we use a batch size of 16, a shared expert rate of 0.75, a constant learning rate of 1e-4, and 400 training steps in merging to obtain FT. Our baseline model, namely SFT, is fine-tuned for 5 epochs with a batch size of 64 and 625 warmup steps for a fair comparison. We also choose the best-performing learning rate from for SFT.
A.10 Training Settings for TinyLlama 1.1B
Using TinyLlama 1.1B as the base model, we upcycle a new MoE model, namely MoE, from the pre-trained dense model. Following the setting of MoE, we construct MoE with 8 experts in one MoE layer, where the top 6 experts are activated for each token, including one shared expert. As such, the number of parameters for MoE can be written as 81.1B. We use a batch size of 64 and a learning rate of 5e-5 with a linear scheduler to fine-tune MoE for 4 epochs with 240 warmup steps. To obtain FT, we learn mixing coefficients to merge MoE layers inside MoE by fine-tuning them with a batch size of 64, a shared expert rate of 0.85, and a learning rate of 2e-5 with a linear schedule for 1 epoch with 60 warmup steps. For a fair comparison, we fine-tune a baseline model SFT for 5 (= 4 + 1) epochs with a batch size of 64, a learning rate of 5e-5, and 300 warmup steps.
A.11 Details of Preliminary Theoretical Explanation
We consider a simplified variant of FT as follows:
@@ A.11 Details of Preliminary Theoretical Explanation
In this simplified scenario, if we denote as the output of the model for the input , the output of this simplified MoE model for input token can be represented as . Interestingly, if we define two new dense models and , where and both use the attention layer of this MoE model as their attention layer while using and as their FFN layer respectively, can be represented as . Consequently, the computation process of this simplified MoE model can be viewed as ensembling the outputs of two dense models and . Meanwhile, the process of merging the upcycled MoE model back to a dense model in this simplified scenario can be represented as , which can be viewed as the model merging of the same two dense models and .